### v1

# EmoSphere++: Emotion-Controllable Zero-Shot Text-to-Speech via Emotion-Adaptive Spherical Vector

Abstract
Emotional text-to-speech (TTS) technology has achieved significant progress in recent years; however, challenges remain owing to the inherent complexity of emotions and limitations of the available emotional speech datasets and models. Previous studies typically relied on limited emotional speech datasets or required extensive manual annotations, restricting their ability to generalize across different speakers and emotional styles. In this paper, we present EmoSphere++, an emotion-controllable zero-shot TTS model that can control emotional style and intensity to resemble natural human speech. We introduce a novel emotion-adaptive spherical vector that models emotional style and intensity without human annotation. Moreover, we propose a multi-level style encoder that can ensure effective generalization for both seen and unseen speakers. We also introduce additional loss functions to enhance the emotion transfer performance for zero-shot scenarios. We employ a conditional flow matching-based decoder to achieve high-quality and expressive emotional TTS in a few sampling steps. Experimental results demonstrate the effectiveness of the proposed framework.
Index Terms:
Emotional speech synthesis, emotion transfer, emotion style and intensity control, zero-shot text-to-speech
publicationid: pubid:
I Introduction
Emotions are interrelated in a highly systematic fashion [1]. For example, the emotion of sadness can be expressed with derivative states of primary emotions such as feeling hurt or lonely, depending on the style and intensity. Recently, emotional text-to-speech (TTS) technology has experienced rapid developments, increasing the interest in high-level interpretable emotion control [2, 3, 4, 5]. Controllable emotional TTS represents a breakthrough in reproducing human-like emotions in speech synthesis, thus enabling more emotionally intelligent interactions between humans and computers. Although researchers have made significant progress in controlling emotional intensity, the ability to precisely control emotional style must be further improved.
The modeling of diverse emotional styles and intensities represents a significant challenge in controllable emotion TTS. A reason for this challenge is that emotion style and intensity are even more subjective and complex than discrete emotion categories, making them difficult to model. Two general approaches for achieving controllable emotional TTS involve controlling conditioning features or manipulating internal emotion representations. That is, one approach uses conditioning features of emotion intensity, such as relative ranking matrices [6, 7, 8, 9, 10, 11], distance-based quantization [12], or voiced, unvoiced, and silence (VUS) states [13]. The alternative approach involves the manipulation of internal emotion representations through the application of scaling factors [14, 15] or interpolation of the embedding space [16]. However, despite these methods, the explicit control of emotion style and intensity remains a largely unexplored topic in emotional speech synthesis.
Figure 1: (a) Three-dimensional valence-arousal-dominance (VAD) cubes of emotions, where all emotional styles occur as derivative states of primary emotions. Emotional intensity control method is used for (b) conventional models and (c) the proposed model with consideration for emotional style.
Another approach to controlling emotional expression involves utilizing emotional dimensions. Compared to the discrete emotion approach, the dimensional approach, such as Russell’s circumplex model, provides a more precise method for capturing the nuances between different emotional states [1]. Recently, studies on TTS systems have attempted to control emotional attributes through the emotion dimension [17, 18]. In one of these studies, a prosody control block is extended by incorporating the continuous space of arousal and valence to allow interpretable emotional prosody control [18]. Another study proposes an expressive TTS model with a semi-supervised latent variable to control emotions in six discrete emotional states of arousal-valence combinations [17]. However, this setup requires labor-intensive annotations, which are more expensive to obtain than categorical labels and more susceptible to annotator bias. The emotional dimension model also exhibits limitations when explicitly controlling emotion style and intensity.
Most emotional TTS systems utilize Sequence-to-sequence (Seq2Seq) models, which not only predict the duration of speech automatically but also learn feature mapping and alignment simultaneously [19, 20]. The attention mechanism in these models allows them to focus on the emotionally emphasized parts of an utterance [21]. However, Seq2Seq models face the typical challenges of auto-regressive models, such as long-term dependence and repetition problems. Furthermore, most emotional speech syntheses adopt fine-tuning to control emotion intensity on a single speaker dataset; however, some of these methods exhibit noticeably degraded speech quality [7, 10]. Researchers have explored acoustic models for emotion transfer to ensure the capture of expressiveness when synthesizing acoustic features. Existing methods focus on emotion transfer using discrete emotion labels, which overlook the complexity of emotions conveyed in human speech [15, 22]. To address this issue, a TTS system with high generalization capability in zero-shot emotion transfer scenarios must be implemented [23, 24, 25, 26, 2].
Figure 2: Illustration of coordinate transformations: (a) Cartesian-to-spherical coordinate transformation and (b) Emotion-adaptive coordinate transformation.
In psychology, several studies [27, 28] have been conducted to explore the frameworks and methods for measuring complex emotions that arise from more primary emotional states. Moreover, TTS models are being challenged by the increasing demand for personalized speech generation, focusing on achieving high generalization capability and producing high-quality speech. However, the following primary research problems exist: (1) how can emotional style and intensity be defined and modeled as derivatives of primary emotions, (2) how can high-quality emotional style be controlled and transferred in zero-shot scenarios, and (3) how can the synthesized emotional speech be evaluated. In this article, we address these challenges, with our primary contributions listed as follows:
  • We introduce an emotion-adaptive coordinate transformation that models an emotion-adaptive spherical vector (EASV), allowing for subtle control over emotion style and intensity.
  • We introduce a multi-level style encoder for high- and low-level style adaptations of custom utterances. Particularly, we introduce additional loss functions to enhance the emotion transfer performance for zero-shot scenarios.
  • We propose a novel objective evaluation of spherical vector angle similarity (SVAS) to assess the emotional expressiveness of the speech.
  • We carefully designed objective and subjective evaluations to demonstrate the effectiveness and contributions of the proposed model from multiple perspectives.
Audio samples and source code are available at https://github.com/Choddeok/EmoSpherepp.
II Background and Related Work
II-A Characterization of Emotions
The processes of defining and expressing emotions have garnered significant interest in psychology [29, 30, 1, 31]. Emotion theorists divide emotion theory into discrete [32, 27] and dimensional models [1, 33]. Notably, discrete models only represent the most straightforward way, whereas dimensional models aim to offer a continuous and fine-grained description.
Emotion labels correspond very closely to the categories that we use in our daily lives. Paul Ekman [32] derived six primary emotions: happiness, anger, disgust, sadness, anxiety, and surprise based on universally recognized facial expressions. However, this approach overlooks the nuanced variations of emotions. Plutchik’s emotion wheel [27] proposes eight primary emotions and suggests that all other emotions arise as derivative states of the primary emotions. Furthermore, varying the intensity of primary emotions on the emotion wheel can produce the wide range of emotions that humans experience. Although people can explicitly express emotions, modeling the relationships between discrete emotional states presents a challenge.
Researchers have introduced dimensional models to computationally interpret the relationships between emotional states. Russell’s circumplex model [1] suggests a two-dimensional circular space that spans the independent and bipolar dimensions of arousal and valence. Building on this, researchers have attempted to extend the model to the third dimension of dominance to denote the location of emotion within this space [33]. In the valence-arousal space, intensity is often equated with arousal; moreover, Reisenzein demonstrated that using the angle and length of the vector in polar coordinates is the only possible option for interpreting the relationships between emotions [28, 34]. Recently, the speech processing domain has seen various studies on emotion recognition that leverage the emotional dimension [35, 36, 37, 38].
Despite these efforts in psychology, the current literature on speech synthesis is still insufficient in effectively modeling and controlling the subtle variations of emotions. Inspired by several psychological theories [27, 28], we believe that the dimensional model allows speech synthesis models to express derived emotions of primary emotions, as shown in Fig. 1 (a). This approach enables researchers to generate, control, and manipulate a wide range of emotions more easily in real-life applications.
II-B Controllable Emotional Speech Synthesis
Recently, speech synthesis models have exhibited significant developments [39, 40, 41]; therefore, controllable emotional speech synthesis research is being aggressively pursued [4, 3, 5]. Recording and annotating emotional datasets covering multiple emotional styles is incredibly challenging. Researchers typically use the following controllable emotional speech synthesis methods to utilize general emotional datasets: 1) emotion label- and 2) reference-based approaches.
The emotion label-based approach aims to properly model conditioning input to reflect the complex nature of emotions. Researchers typically model emotion intensity using a learned ranking function [42], as employed in [6, 7, 10, 8, 9, 11]. The ranking function [42] seeks a ranking matrix based on the relationships between the low-level and different high-level emotional expressions using support vector machines. The model receives the emotional intensity of emotional samples as a conditioning input for training. However, this method tends to rely on emotion labels and introduces bias into training through separate stages. Most models that use ranking functions adopt fine-tuning to control emotion intensity on a single-speaker dataset; however, some of these methods have noticeably degraded speech quality [7, 10]. Moreover, certain research studies have utilized conditioning input such as distance-based quantization [12] and VUS states [13] to model emotional intensity. However, these methods are still limited to several predefined emotion labels and lack differentiation among samples within the same emotion label.
As emotional speech synthesis often lacks multiple emotional style labels, reference-based approaches are famous for using reference audio to transfer emotional styles. Several studies have controlled emotion intensity through operations on representative emotion embedding. The scaling factors approach [14, 15] reflects fine-grained emotion representation through multiplication. In addition to the scaling approach, the interpolation approach proposed by [16] controls emotion intensity through an inter-to-intra emotional distance ratio algorithm. However, the structure of the embedding space influences model performance and complicates the process of finding proper parameters for scaling or interpolation.
However, these frameworks can neither be explicitly tuned such as label-based approaches nor capture fine-grained emotion representations such as reference-based approaches. This study aims to fill this research gap by explicitly modeling EASV of emotion style and intensity variations.
II-C Style and Emotion Transfer in Text-to-Speech
Researchers in the TTS community have studied style and emotion transfer for decades. Some studies introduced the concept of global style tokens (GSTs) [43]. One such derivative is Mellotron [44], an autoregressive multi-speaker TTS model based on Tacotron that utilizes GSTs. Li et al. [15] proposed a module for disentangling and transferring emotions across different speakers to achieve a desired emotional tone while maintaining the identity of the speaker. However, these methods exhibit limitations in capturing different styles and zero-shot scenarios owing to the restricted speaker lookup table and the insufficient performance of the disentangling modules.
Conversely, iEmoTTS [45] and YourTTS [46] used pre-trained speaker embedding for robust zero-shot performance. Moreover, GenerSpeech [47] proposed a multi-level style adapter to obtain different styles, including a global latent representation with speaker and emotion features. However, previous research has lacked methods to effectively process style and disentangle speech factors. Our work handles well-formed style processing through multi-level style encoders and incorporation of additional loss to exhibit strong generalization performance for zero-shot scenarios.
Figure 3: Training diagram of the EmoSphere++ framework. The framework consists of three main modules: the text encoder, the multi-level style encoder, and the CFM decoder. The right section illustrates the detailed structure of the multi-level style encoder, which extracts high-level speaker, high-level emotion, and low-level emotion to form a multi-level style embedding for emotional speech synthesis.
III Proposed Method
This paper introduces EmoSphere++, an emotion-controllable zero-shot TTS model that can control emotional style and intensity to resemble natural human speech. We considered the emotion wheel theory [27], which suggests that all other emotions are derived from the primary emotions, and the circumplex model [1, 34], which represents emotions through coordinate transformations based on intensity. Inspired by these psychological theories, we believe that derivative emotions can be characterized by the coordinate transformation in the valence-arousal-dominance (VAD) dimension, reflecting the style and intensity of the primary emotions, as shown in Fig. 1 (c). In a previous work, a spherical emotion vector was introduced to model the style and intensity of emotions [3]. Building on this, we propose a novel EASV, enabling more precise manual manipulation of the synthesized emotional style and intensity. This framework facilitates the training of each high-level style pre-trained encoder with disentangling modules, allowing us to effectively generalize across zero-shot scenarios. The details are described in the following subsections.
III-A Emotion-Adaptive Coordinate Transformation
As Fig. 2 (a) shows, the simplest way to determine the intensity of an emotion $k$ is to assume that the emotion intensity decreases as it moves closer to the neutral emotion center $M$ as follows [3, 12]:
$M=\frac{1}{N_{n}}\sum_{i=1}^{N_{n}}e_{i}^{n}.$ (1)
Here, $N_{n}$ is the total number of a neutral coordinate $e_{ni}$. However, as the mean-based method does not consider the distribution of other emotion categories, such as variance, it does not fully capture the information of the target emotion. In our proposal, we simultaneously consider the distribution of the target and other emotion categories and extract an adaptive spherical vector of each emotion category.
We studied a coordinate method to measure the angle-based distance measure, which can model emotional style and intensity on a continuous scale [3, 34]. The key idea of the proposed method is to model a spherical coordinate system for each emotion by simultaneously considering the distribution of the target and other emotion categories, as shown in Fig. 2 (b). We first considered two assumptions: (1) the emotional intensity increases as it moves farther from the center of the emotion-adaptive spherical coordinate system and (2) the angle from the center of the emotion-adaptive spherical coordinate determines the emotional style. Initially, we adopted a specific speech emotion recognition (SER) model [48] $\psi$ to predict the VAD value $e_{i}^{k}$ in emotion class $k$:
$e_{i}^{k}=\psi(x_{i}),$ (2)
where $x_{i}$ denote the $i$-th referece speech in speech dataset $\mathrm{X}$ and $e_{i}^{k}$ consists of values for $(d_{v},d_{a},d_{d})$, where $d_{v}$, $d_{a}$, and $d_{d}$ represent valence, arousal, and dominance, respectively. Each component is expressed in Cartesian coordinates, with values ranging from 0 to 1. To model the spherical coordinate system for each emotion, we obtained the shifted Cartesian coordinates $\widehat{e_{i}^{k}}=(\widehat{d}_{v},\widehat{d}_{a},\widehat{d}_{d})$ by shifting through the representative central coordinates $M_{k}$ from different emotion coordinate set $E_{k}$. The coordinates $M_{k}$ are extracted using emotion-specific centroid extraction, which maximizes the ratio of the distance between specific emotions to the distance from the neutral coordinates as follows:
Input: Emotional speech dataset $\mathrm{X}$, SER model $\psi$
1
Output: Emotion-adaptive spherical vector set $\mathbb{S}$
2
3for (reference speech $x_{i}$, emotion class $k$) in $\mathrm{X}$ do:
4 Get VAD value $e_{i}^{k}\in E_{k}$ using SER model $\psi$ by Eq. (2)
5end for
6for each emotion class $k$ of $E_{k}$ do:
7if $E_{k}$ is the neutral class set $E_{n}$ then:
8  for $e_{i}^{n}$ in $E_{n}$ do:
9   Append $s_{i}^{n}=(0,0,0)$ to $\mathbb{S}$
10  end for
11else:
12  Compute centroid coordinate as $M_{k}$ by Eq. (3)
13  for $e_{i}^{k}$ in $E_{k}$ do:
14   Compute shifted VAD as $\widehat{e_{i}^{k}}$ by Eq. (4)
15   Spherical transformation as $s_{i}^{k}\in S_{k}$ by Eq. (5)
16  end for
17  Calculate ($r_{min}$, $r_{max}$) the interquartile range of $r$
18  for $s_{i}^{k}$ in $S_{k}$ do:
19   Compute $r_{IQR}$ by Eq. (6)
20   Append $s_{i}^{k}=(r_{IQR},\vartheta,\varphi)$ to $\mathbb{S}$
21  end for
22end for
Algorithm 1 Emotion-Adaptive Coordinate Transformation
$M_{k}=\arg\max_{M}\frac{\mathbb{E}_{e_{i}^{k}\in E_{k}}[\|M-e_{i}^{k}\|_{2}]}{\mathbb{E}_{e_{i}^{n}\in E_{n}}[\|M-e_{i}^{n}\|_{2}]},$ (3)
$\widehat{e_{i}^{k}}=e_{i}^{k}-M_{k}.$ (4)
Here $e_{i}^{k}$ and $e_{i}^{n}$ denote the $i$-th coordinate of the $k$-th emotion coordinate set $E_{k}$ and the neutral coordinate set $E_{n}$, respectively. Consequently, the centroid coordinates are the ones that maximize the distances from the target emotion category while minimizing the distance within the neutral emotion category. Then, transformation via the representative central coordinates $M_{k}$ to spherical coordinates $s_{i}^{k}=(r,\vartheta,\varphi)$ can be formulated as follows:
$r=\sqrt{{\widehat{d_{v}}}^{2}+{\widehat{d_{a}}}^{2}+{\widehat{d_{d}}}^{2}},$
$\vartheta=\arccos\left(\frac{\widehat{d}_{d}}{r}\right),\varphi=\arctan\left(\frac{\widehat{d}_{v}}{\widehat{d}_{a}}\right).$ (5)
After the emotion-adaptive coordinate transformation, we applied the interquartile range (IQR) technique [49] to adjust the data based on the median, thereby reducing the influence of outliers as follows:
$r_{clamp}=\min(\max(r,r_{min}),r_{max}),$
$r_{IQR}=\frac{r_{clamp}-r_{min}}{r_{max}-r_{min}}.$ (6)
Here $r_{min}$ and $r_{max}$ are bounds derived based on the IQR technique with $r_{min}$ set as the first quartile minus 1.5 times the IQR and $r_{max}$ as the third quartile plus 1.5 times the IQR. The detailed procedure for obtaining the emotion-adaptive spherical vector set $\mathbb{S}$ is outlined in Algorithm 1.
III-B Multi-Level Style Encoder
The multi-level style encoder consists of two main modules: the emotion and speaker encoders. The emotion encoder comprises high- and low-level emotion encoders.
As shown in Fig. 3, the emotion encoder includes a fine-tuned SER model for high-level emotion features and an additional EASV extractor for low-level emotion features. The high-level emotion encoder extracts the fixed-size hidden embedding from the SER [^1] https://github.com/ddlBoJack/emotion2vec [50], which contains a multilayer transformer. Meanwhile, the EASV extractor, which is described in Section III-A, generates low-level emotion features. A fully connected layer processes high- and low-level emotion features, combining them into a fixed-size hidden embedding.
The speaker encoder provides speaker-related information to the TTS model. A pretrained speech encoder extracts a speaker embedding [51] from the reference speech for zero-shot emotion transfer. We used the WavLM Base model [52] as the speaker verification model [^2] https://huggingface.co/microsoft/wavlm-base-sv, building the model on the HuBERT [53] framework to focus on modeling speech content and preserving speaker identity. WavLM-based approaches can capture and represent speaker-specific information and emotional nuances in speech [54, 55]. Similar to emotion embeddings, fixed-size hidden embeddings are processed in the fully connected layer and combined to generate multi-level style embedding $e_{sty}$.
III-C Preliminary on Conditional Flow Matching-Based Model
EmoSphere++ adopts a conditional flow matching (CFM)-based decoder that generates flow through the ordinary differential equation. Building on the success of flow matching in the speech synthesis task [23, 25, 24], we utilized a CFM-based decoder that is designed to model a conditional vector field $\mathbf{u}_{t}$. Following [56], we define the flow $\phi$ as the mapping between two density functions:
$\tfrac{d}{dt}\phi_{t}(x)=\boldsymbol{v}_{t}(\phi_{t}(x))\text{;}\quad\quad\phi_{0}(x)=x\text{.}$ (7)
Here, $\boldsymbol{v}_{t}$ represents a time-dependent vector field that defines the path of the probability flow over time $t\in\left[0,1\right]$. Particularly, it describes a conditional flow process in which the conditional flow $\phi_{t,x1}$ represents simple linear trajectories between the data point $x_{1}$ drawn from the target distribution $q(x)$ and prior distribution $x_{0}\sim N(0,I)$:
$\phi_{t,x1}(x_{0})=(1-(1-\sigma_{min})t)x_{0}+tx_{1},$ (8)
where $\sigma_{min}$ is the hyper-parameter for small amounts of white noise. The vector field in the decoder is trained using the following objectives:
$\mathit{L}_{cfm}=\mathbb{E}_{t,q(x_{1}),p(x_{0})}\left\|u_{t}(\phi_{t,x_{1}}(x_{0}))-\widetilde{\upsilon}_{\theta}(\phi_{t,x_{1}}(x_{0}),\mu,e_{sty},t)\right\|^{2},$ (8)
where $\mu$ represents the predicted average acoustic features (e.g., Mel-spectrogram) given the text and the chosen durations, using a text encoder and duration predictor. $e_{sty}$ denotes the multi-level style embedding.
III-D Training Objective
Alongside traditional losses for training TTS systems, we introduce objective functions to help in modeling our system to control and transfer emotion style and intensity.
First, text encoder and duration predictor architectures were implemented based on [23]. Duration-model training uses monotonic alignment search to compute the duration loss $L_{dur}$ and the prior loss $L_{enc}$, as described in [57]. The decoder follows the CFM loss $L_{cfm}$ described in Section III-C.
Inspired by [15], we introduced an additional orthogonality loss. In the style encoder, emotion and speaker embeddings contain overlapping information regarding each other. The speaker and emotion embeddings should be 1) discriminative in distinguishing identities and 2) independent of each other, ensuring effective generalization performance for both seen and unseen speakers. To mitigate the impact of speaker and emotion embedding leakage on model performance, we propose a normalized orthogonality loss $L_{ort}$ to enhance the decoupling capability of the model. Unlike the existing loss, this method normalizes all sample pairs, providing effective generalized performance for zero-shot scenarios. In this case, $L_{ort}$ can be expressed as:
$L_{ort}=\sum_{j=1}^{n}\sum_{i=1}^{n}\frac{\left\|s_{i}^{T}e_{j}\right\|^{2}}{\left\|s_{i}\right\|^{2}\left\|e_{j}\right\|^{2}},$ (9)
[…]
Figure 4: Run-time diagram of the proposed EmoSphere++ framework. The emotion-adaptive spherical vector can manually control the low-level emotion of emotion style and intensity. We produce an emotional state as the derivative of primary emotions by assigning the appropriate angle and length to the spherical vector.
Consequently, the final objective function is defined as:
$L_{total}=\lambda_{enc}L_{enc}+\lambda_{cfm}L_{cfm}+\lambda_{dur}L_{dur}+\lambda_{ort}\mathcal{L}_{ort},$ (10)
where $\lambda_{enc}$, $\lambda_{cfm}$, $\lambda_{dur}$, and $\lambda_{ort}$ are the loss weights, which we set to 1.0, 1.0, 1.0, and 0.02, respectively.
III-E Control of Emotional Style and Intensity
Fig. 4 illustrates the proposed emotion-controllable zero-shot TTS framework, which synthesizes emotional speech for both seen and unseen speakers based on reference speech and EASV. The framework comprises three main modules: the text encoder, multi-level style encoder, and CFM decoder.
The multi-level style encoder captures the emotion and speaker information in an embedding from the reference speech. By varying the length and angle in the EASV, we can manipulate the levels of style and intensity at runtime and efficiently synthesize the desired emotional effects. In a zero-shot scenario, the framework can synthesize speech for unseen speakers by inputting the unseen target reference speech into the speaker encoder.
TABLE I: Prosodic variation analysis of emotion-adaptive spherical vector values on ESD datasets. Q1, Q2, and Q3 represent intensity regions divided by 0.33 and 0.66 thresholds. The average prosodic change $p_{c}$ shows the ratio of the average of each style to the overall emotion average. “-” indicates missing data, with neutral emotions fixed at intensity 0. Cell hue represents heat-mapped prosody values for each emotion.
Emotion Style Num. Pitch (mean) Energy (mean) Duration (mean)
Q1 Q2 Q3 All Q1 Q2 Q3 Avg. $p_{c}$ Q1 Q2 Q3 Avg. $p_{c}$ Q1 Q2 Q3 Avg. $p_{c}$
Neutral All (Average) - - - 3,500 - - - 51.68 - - - - 0.41 - - - - 2.60 -
Angry I (+V +A +D) 297 788 137 1,222 55.00 66.02 76.79 64.68 +0.14 0.99 1.43 1.86 1.37 +0.69 2.96 2.85 2.81 2.87 +0.05
IV (+V -A +D) 1 - - 1 47.78 - - 47.78 -0.16 0.67 - - 0.67 -0.18 2.28 - - 2.28 -0.17
V (+V +A -D) 374 1,365 533 2,272 53.87 65.64 78.65 66.92 +0.18 1.05 1.37 1.73 1.40 +0.73 2.79 2.76 2.72 2.76 +0.01
VII (-V -A -D) 3 - - 3 51.96 - - 51.96 -0.08 0.97 - - 0.97 +0.20 2.85 - - 2.85 +0.04
VIII (+V -A -D) 2 - - 2 57.29 - - 57.29 +0.01 0.02 - - 0.02 -0.98 2.65 - - 2.65 -0.03
All (Average) 677 2,153 670 3,500 54.33 65.78 78.26 56.66 - 1.01 1.40 1.76 0.81 - 2.87 2.79 2.74 2.75 -
Happy I (+V +A +D) 338 1,482 472 2,292 64.46 77.01 90.05 77.84 +0.62 0.59 0.83 0.99 0.83 +0.83 2.77 2.67 2.66 2.68 -0.11
II (-V +A +D) 4 5 - 9 35.165 33.56 - 34.24 -0.29 0.23 0.16 - 0.19 -0.58 2.94 3.31 - 3.15 +0.04
III (-V -A +D) 81 18 - 99 31.99 28.20 - 31.37 -0.35 0.21 0.17 - 0.20 -0.56 3.38 3.38 - 3.38 +0.12
IV (+V -A +D) 20 - - 20 52.54 - - 52.54 +0.09 0.39 - - 0.39 -0.13 2.96 - - 2.96 -0.02
V (+V +A -D) 159 666 171 996 62.93 78.66 89.26 78.09 +0.63 0.67 0.90 1.34 0.94 +1.08 2.69 2.58 2.66 2.61 -0.13
VI (-V +A -D) 2 - - 2 31.01 - - 31.01 -0.36 0.08 - - 0.08 -0.83 3.60 - - 3.60 +0.19
VII (-V -A -D) 56 12 - 68 32.28 26.74 - 31.44 -0.35 0.22 0.23 - 0.22 -0.51 3.36 3.07 - 3.31 +0.10
VIII (+V -A -D) 14 - - 14 47.75 - - 47.75 -0.01 0.76 - - 0.76 +0.68 2.41 - - 2.41 -0.19
All (Average) 674 2,183 643 3,500 55.16 76.63 89.84 48.03 - 1.31 1.97 2.51 0.45 - 2.88 2.65 2.66 3.02 -
Sad I (+V +A +D) 13 29 1 43 61.86 76.17 90.43 72.17 +0.13 0.94 1.04 2.00 1.04 +0.27 2.82 2.71 2.92 2.75 -0.05
II (-V +A +D) 1 5 - 6 33.82 54.10 - 51.03 -0.21 0.89 0.65 - 0.70 -0.14 3.89 3.03 - 3.17 +0.1
III (-V -A +D) 184 713 98 995 57.95 53.21 43.51 53.15 -0.17 0.72 0.60 0.38 0.60 -0.27 2.95 3.06 3.25 3.06 +0.06
IV (+V -A +D) 39 11 - 50 72.76 75.94 - 73.42 +0.14 0.88 0.95 - 0.89 +0.09 2.83 2.68 - 2.79 -0.03
V (+V +A -D) 46 113 15 174 68.97 76.37 67.01 73.68 +0.15 1.11 1.03 0.87 1.04 +0.28 2.97 2.78 2.31 2.79 -0.03
VI (-V +A -D) 3 4 - 7 55.81 65.45 - 62.01 -0.03 0.67 0.67 - 0.67 -0.18 2.60 3.02 - 2.84 -0.02
VII (-V -A -D) 228 1,336 535 2,099 62.74 53.83 38.81 51.10 -0.20 0.70 0.67 0.52 0.64 -0.22 2.91 2.93 3.15 2.98 +0.03
VIII (+V -A -D) 54 70 2 126 74.43 78.60 82.04 76.84 +0.20 0.99 0.93 0.88 0.96 +0.17 2.67 2.71 2.98 2.70 -0.07
All (Average) 568 2,281 651 3,500 63.36 55.87 40.28 64.18 - 0.79 0.68 0.51 0.82 - 2.89 2.95 3.15 2.89 -
Surprise I (+V +A +D) 339 1,083 279 1,693 61.13 83.05 93.42 80.48 +0.35 0.98 1.33 1.88 1.35 +0.15 2.87 2.75 2.76 2.78 -0.01
IV (+V -A +D) 3 - - 3 46.2 - - 46.2 -0.23 0.75 - - 0.75 -0.36 3.5 - - 3.5 +0.25
V (+V +A -D) 322 1,079 385 1,786 60.68 82.14 103.48 82.89 +0.39 1.01 1.33 2.09 1.43 +0.22 2.77 2.7 2.56 2.69 -0.04
VII (-V -A -D) 1 - - 1 48.73 - - 48.73 -0.19 1.35 - - 1.35 +0.16 2.49 - - 2.49 -0.11
VIII (+V -A -D) 9 - - 9 50.14 - - 50.14 -0.16 1.26 - - 1.26 +0.08 2.53 - - 2.53 -0.1
All (Average) 674 2,162 664 3,500 60.59 82.60 99.16 59.77 - 0.99 1.33 2.00 1.17 - 2.82 2.73 2.65 2.80 -
IV Experiments
IV-A Experimental Setup
We conducted experiments using the emotional speech dataset (ESD) [^3] https://github.com/HLTSingapore/Emotional-Speech-Data[58], which contains 350 parallel utterances spoken by 10 English speakers in five emotional states (neutral, happy, angry, sad, and surprise). Following the prescribed data partitioning criteria, we extracted one sample for each emotion from every speaker, resulting in 17,500 samples. The validation set comprised 20 samples for each emotion per speaker, totaling 1,000 samples, whereas the test set comprised 30 samples for each emotion per speaker, totaling 1,500 samples. The zero-shot scenario used two unseen speakers, one English-speaking male (“0013”) and one English-speaking female (“0019”), by excluding them from the training process.
Moreover, we conducted experiments on the interactive emotional dyadic motion capture database (IEMOCAP) [59] to verify whether the models can reflect the styles using predicted VAD values. The IEMOCAP contains 10 speakers with nine emotions (angry, excited, fear, sad, surprised, frustrated, happy, disappointed, and neutral) and dimensional labels such as VAD. The validation set included 10 samples for each speaker, whereas the test set comprised 15 samples.
For the Mel-spectrogram, we transformed audio using the short-time Fourier transform with a hop size of 256, a window size of 1,024, an fast Fourier transform size of 1,024, and 80 Mel bins. We converted the text to phoneme using the grapheme-to-phoneme tool of the Festival Speech Synthesis System [60] to serve as the input to the text encoder. We designed parallel and non-parallel test scenarios at runtime depending on whether the input text matches the reference speech.
TABLE II: Subjective and objective evaluation results for non-parallel emotion transfer on the seen dataset.
The nMOS, sMOS, and eMOS scores are presented with 95% confidence intervals.
Method Subjective Evaluation Objective Evaluation
nMOS ($\uparrow$) sMOS ($\uparrow$) eMOS ($\uparrow$) WERWhis ($\downarrow$) WERw2v ($\downarrow$) WERAVG ($\downarrow$) SECSR ($\uparrow$) SECSW ($\uparrow$) SECSAVG ($\uparrow$) ECA ($\uparrow$) SVAS ($\uparrow$) EECS ($\uparrow$)
GT 4.06$\pm$0.05 4.15$\pm$0.05 4.11$\pm$0.05 11.42 14.99 13.21 0.7358 0.9264 0.8311 95.53 - 0.9487
BigVGAN [61] 3.95$\pm$0.05 4.01$\pm$0.06 3.98$\pm$0.06 11.36 15.15 13.26 0.7271 0.9231 0.8251 94.25 0.9815 0.9389
Mellotron [44] 3.42$\pm$0.07 3.95$\pm$0.07 3.75$\pm$0.07 14.49 18.90 16.70 0.6981 0.8939 0.7960 46.93 0.7838 0.5061
Mixedemotion [10] 3.36$\pm$0.07 3.89$\pm$0.07 3.69$\pm$0.07 20.24 26.48 23.36 0.6920 0.8764 0.7842 63.44 0.8669 0.6282
YourTTS [46] 3.52$\pm$0.06 3.94$\pm$0.06 3.75$\pm$0.07 28.71 32.79 30.75 0.7296 0.6751 0.6811 74.84 0.8657 0.7890
Generspeech [47] 3.85$\pm$0.06 3.96$\pm$0.06 3.86$\pm$0.06 16.63 22.56 19.60 0.7074 0.8888 0.7981 82.54 0.8460 0.8366
iEmoTTS [45] 3.77$\pm$0.06 3.94$\pm$0.06 3.79$\pm$0.07 26.07 29.74 27.91 0.6153 0.8208 0.7181 77.60 0.8000 0.7474
EmoSphere ++ (Proposed) 3.92$\pm$0.06 3.97$\pm$0.06 3.86$\pm$0.06 15.52 18.85 17.19 0.7314 0.9047 0.8181 93.53 0.8717 0.9270
TABLE III: Subjective and objective evaluation results for non-parallel emotion transfer on the unseen dataset.
The nMOS, sMOS, and eMOS scores are presented with 95% confidence intervals.
Method Subjective Evaluation Objective Evaluation
nMOS ($\uparrow$) sMOS ($\uparrow$) eMOS ($\uparrow$) WERWhis ($\downarrow$) WERw2v ($\downarrow$) WERAVG ($\downarrow$) SECSR ($\uparrow$) SECSW ($\uparrow$) SECSAVG ($\uparrow$) ECA ($\uparrow$) SVAS ($\uparrow$) EECS ($\uparrow$)
GT 4.02$\pm$0.05 4.04$\pm$0.05 4.08$\pm$0.05 13.06 15.15 14.11 0.7725 0.9282 0.8504 100.00 - 0.9944
BigVGAN [61] 3.93$\pm$0.06 4.01$\pm$0.05 4.04$\pm$0.06 13.11 17.07 15.09 0.7603 0.9275 0.8439 99.32 0.9844 0.9903
YourTTS [46] 3.67$\pm$0.07 3.93$\pm$0.05 3.71$\pm$0.08 26.48 32.78 29.63 0.6870 0.8220 0.7545 61.46 0.8323 0.6509
Generspeech [47] 3.90$\pm$0.06 3.96$\pm$0.06 3.83$\pm$0.07 15.42 23.46 19.44 0.6812 0.8191 0.7502 80.57 0.8131 0.7935
iEmoTTS [45] 3.83$\pm$0.06 3.93$\pm$0.05 3.82$\pm$0.08 26.74 32.70 29.72 0.5970 0.7480 0.6725 51.03 0.7615 0.5364
EmoSphere ++ (Proposed) 3.91$\pm$0.06 3.99$\pm$0.05 3.85$\pm$0.07 14.41 18.43 16.42 0.6543 0.8640 0.7592 94.61 0.8458 0.9358
IV-B Implementation Details
For the acoustic model, we followed the Matcha-TTS [23] configuration, utilizing the text encoder and duration predictor in the encoder, along with a 1D U-Net based CFM decoder. The CFM decoder consists of two downsampling blocks, followed by two midblock and two upsampling blocks, each containing a Transformer layer with a hidden dimensionality of 256, an attention module with dimensionality of 64, and “snakebeta” activations. Following [62], the text encoder was modified by incorporating relative position representations and adding a residual connection to the encoder pre-net. Based on [63], the duration predictor consists of two convolutional layers with rectified linear unit activation, followed by layer normalization, dropout, and a projection layer. Regarding the SER444huggingface.co/wav2vec2-large-robust-12-ft-emotion-msp-dim, we adopt a system proposed in [48], which predicts VAD using wav2vec 2.0 [64] and a linear predictor. We trained the generator using random segments of 32 frames from the Mel-spectrogram, with a batch size of 32 and a total of 11M training steps. The AdamW optimizer was used, with a $1\times 10^{-4}$ learning rate. In the inference stage, the guidance level $\gamma$ was set to 100. We trained the vocoder using the official BigVGAN [^5] https://github.com/NVIDIA/BigVGAN[61] implementation, incorporating LibriTTS [65], voice cloning toolkit [^6] https://datashare.ed.ac.uk/handle/10283/2651, and ESD datasets. All comparison models were trained using a single NVIDIA RTX A6000 GPU.
IV-C Evaluation
IV-C1 Subjective Metrics
We adopted two subjective metrics: 1) a mean opinion score (MOS) evaluation and 2) a preference test to evaluate emotion expressiveness. We conducted subjective evaluations using Amazon Mechanical Turk. All subjects were required to listen with headphones and replay each sample 2-3 times.
We conducted a MOS evaluation for naturalness (nMOS), speaker similarity (sMOS), and emotion similarity (eMOS) using a nine-point scale ranging from 1 to 5, with increments of 0.5 units. The results are presented with a confidence interval of 95%. We randomly selected 10 utterances from the evaluation set.
We also conducted a preference test to evaluate the modeling of emotion expressiveness. To demonstrate the success of our modeling, we synthesized speech with three different levels of emotion intensity (weak, medium, and strong) from pairs of speech that share the same emotion and style. Evaluators were presented with two different sentences with the same emotion and style, each with varying intensities, and tasked with selecting the one exhibiting a stronger emotion. We uniformly referred to scores 0.1 as weak, 0.5 as medium, and 0.9 as strong and randomly choose 10 utterances from the evaluation set.
IV-C2 Objective Metrics
[…]
TABLE IV: Comparison with the results of control methods for non-parallel emotion transfer on the seen dataset.
The nMOS, sMOS, and eMOS scores are presented with 95% confidence intervals.
Method Subjective Evaluation Objective Evaluation
nMOS ($\uparrow$) sMOS ($\uparrow$) eMOS ($\uparrow$) WERWhis ($\downarrow$) WERw2v ($\downarrow$) WERAVG ($\downarrow$) SECSR ($\uparrow$) SECSW ($\uparrow$) SECSAVG ($\uparrow$) ECA ($\uparrow$) SVAS ($\uparrow$) EECS ($\uparrow$)
GT 4.06$\pm$0.05 4.15$\pm$0.05 4.11$\pm$0.05 11.42 14.99 13.21 0.7358 0.9264 0.8311 95.53 - 0.9487
BigVGAN [61] 3.95$\pm$0.05 4.01$\pm$0.06 3.98$\pm$0.06 11.36 15.15 13.26 0.7271 0.9231 0.8251 94.25 0.9815 0.9389
MatchaTTS w/ Scaling Factor [15] 3.81$\pm$0.06 3.95$\pm$0.06 3.82$\pm$0.07 18.36 23.57 20.97 0.7212 0.9012 0.8112 40.63 0.8191 0.4757
MatchaTTS w/ Relative Attribute [7] 3.86$\pm$0.06 3.94$\pm$0.06 3.84$\pm$0.07 15.53 19.87 17.70 0.7335 0.9066 0.8201 85.31 0.8672 0.8756
EmoSphere ++ (Proposed) 3.92$\pm$0.06 3.97$\pm$0.06 3.86$\pm$0.06 15.52 18.85 17.19 0.7314 0.9047 0.8181 93.53 0.8717 0.9270
Figure 5: Pitch tendency track according to intensity for different emotions. Pitch values were calculated by averaging the synthesized speech for each intensity across all test sentences.
Figure 6: Pitch tendency track according to intensity for an unseen speaker.
Moreover, we propose the SVAS to evaluate the emotion of synthesized speech. The SVAS is obtained by computing the cosine similarity of the angle of the emotion spherical vector between the synthesized audio and reference audio with the target emotion. We used a SER model101010huggingface.co/wav2vec2-large-robust-12-ft-emotion-msp-dim [48] to predict the VAD values and applied a Cartesian-to-spherical transformation through fixed neutral center coordinates to extract the angle of the emotion spherical vector.
IV-D Comparison Models
We compared the similarity and quality of samples generated using the proposed EmoSphere++ with those produced by other systems. We used the same vocoder and open code, and the comparative models are summarized as follows:
  • GT and BigVGAN [61]: The ground truth audio and waveforms are generated from the ground truth Mel-spectrogram using vocoder.
  • Mellotron [44]: This auto-regressive multi-speaker TTS model for direct style control by conditioning on rhythm and pitch utilizes conditioning GSTs.
  • Mixedemotion [10]: It is a relative attribute ranking-based model pre-computes intensity values for mixed emotion synthesis and allows manual control at run time.
  • YourTTS [46]: It is a zero-shot multi-speaker TTS with the pre-trained speaker encoder. Unlike other comparison models, it does not use a pre-trained vocoder through end-to-end training.
  • GenerSpeech [47]: It is a high-fidelity zero-shot style transfer method for out-of-domain TTS. We used the emotion2vec model instead of the fine-tuned wav2vec 2.0 model to capture global style for a fair comparison.
  • iEmoTTS [45]: This non-autoregressive TTS model for a cross-speaker emotion transfer system has been developed based on timbre-prosody disentanglement.
  • EmoSphere++: This is the proposed CFM-based emotion-controllable zero-shot TTS with EASV.
To ensure a fair comparison between the proposed method and existing emotion intensity modeling approaches, we trained the models using the same dataset and configurations as Matcha-TTS [23], including the encoder (i.e., text encoder and duration predictor) and the U-Net-based decoder. We replaced the speaker and emotion attribute control module for a contrastive study with two competing controllable emotion methods: scaling factor and relative attributes through comprehensive experiments.
  • Matcha-TTS w/ Scaling Factor: Here, the emotion embedding, extracted through the emotion disentangling module, is multiplied by a scaling factor to control the emotion strength [15]. The system manages the speaker identity controller using speaker embeddings obtained from a speaker look-up table.
  • Matcha-TTS w/ Relative Attributes: Here, the relative attributes vector is obtained from a ranking function to control the emotion strength [7]. We employed the same emotion encoder and speaker look-up table, except for fine-tuning the emotion encoder with a single-speaker emotion dataset.
In summary, we conducted emotion transfer experiments to compare EmoSphere++ with other systems, including Mellotron, Mixedemotion, YourTTS, GenerSpeech, and iEmoTTS. Moreover, we evaluated emotion intensity control by comparing it with other intensity control methods, such as Matcha-TTS w/ Scaling Factor and Matcha-TTS w/ Relative Attributes.
Figure 7: Pitch tracks of a sample demonstrating the effects of emotional style shift in sad emotion, where $A$, $V$, and $D$ represent arousal, valence, and dominance, respectively. The line color represents emotional intensity: red = 0.1, green = 0.5, and blue = 0.9.
Figure 8: Comparison of the emotion classification accuracy scores of shifting emotion style.
V Results
V-A Analysis of Emotional Style and Intensity Modeling
As previously mentioned, we characterized the derivative states of emotions through emotion style and intensity. Table I lists the results of the analysis of the novel approach to modeling emotion style and intensity using EASV. In this study, we divided the emotion space into eight regions (“I”$\sim$“VIII”) based on the VAD axes and then shifted the style along spherical spaces. Here, valence represents the positivity or negativity of an emotion, arousal indicates the intensity of the emotion provoked by a stimulus, and dominance denotes the level of control exerted by the stimulus. We aimed to demonstrate the modeling of diverse emotions by analyzing average prosodic variation rates and prosodic variation with intensity based on the characterization of the VAD axis.
TABLE V: Ablation study on disentangling modules for non-parallel emotion transfer on the seen and unseen datasets.
The nMOS, sMOS, and eMOS scores are presented with 95% confidence intervals.
Method Seen Speaker Unseen Speaker
Subjective Evaluation Objective Evaluation Subjective Evaluation Objective Evaluation
nMOS ($\uparrow$) sMOS ($\uparrow$) eMOS ($\uparrow$) SECSR ($\uparrow$) SECSW ($\uparrow$) SECSAVG ($\uparrow$) ECA ($\uparrow$) EECS ($\uparrow$) nMOS ($\uparrow$) sMOS ($\uparrow$) eMOS ($\uparrow$) SECSR ($\uparrow$) SECSW ($\uparrow$) SECSAVG ($\uparrow$) ECA ($\uparrow$) EECS ($\uparrow$)
GT 4.06$\pm$0.06 4.15$\pm$0.05 4.11$\pm$0.05 0.7358 0.9264 0.8311 95.53 0.9487 4.02$\pm$0.05 4.04$\pm$0.05 4.08$\pm$0.06 0.7725 0.9282 0.8504 100.00 0.9944
BigVGAN 3.95$\pm$0.05 4.01$\pm$0.06 3.98$\pm$0.06 0.7271 0.9231 0.8251 94.25 0.9389 3.93$\pm$0.06 4.01$\pm$0.05 4.04$\pm$0.06 0.7603 0.9275 0.8439 99.32 0.9903
w/o Disentangling Module 3.88$\pm$0.06 3.97$\pm$0.06 3.82$\pm$0.07 0.7289 0.9052 0.8171 92.86 0.9218 3.88$\pm$0.06 3.93$\pm$0.06 3.82$\pm$0.07 0.6430 0.8715 0.7573 88.56 0.8802
w/ Orthogonality Loss [15] 3.86$\pm$0.06 3.95$\pm$0.06 3.83$\pm$0.06 0.7287 0.9032 0.8160 92.26 0.9191 3.92$\pm$0.06 3.93$\pm$0.06 3.81$\pm$0.07 0.6454 0.8644 0.7549 91.23 0.9101
w/ Vector Quantization [47] 3.91$\pm$0.06 3.96$\pm$0.06 3.80$\pm$0.07 0.7258 0.9027 0.8143 92.34 0.9235 3.85$\pm$0.06 3.96$\pm$0.06 3.82$\pm$0.08 0.6170 0.8354 0.4177 92.60 0.9320
EmoSphere ++ (Proposed) 3.92$\pm$0.06 3.97$\pm$0.06 3.86$\pm$0.06 0.7314 0.9047 0.8181 93.53 0.9270 3.92$\pm$0.06 3.99$\pm$0.05 3.85$\pm$0.07 0.6543 0.8640 0.7592 94.61 0.9385
V-A1 Prosodic Variation with Intensity Based on the Valence-Arousal-Dominance (VAD) Axis
In this section, we describe the analysis of how prosodic features, such as pitch, vary with emotion intensity based on the VAD axis. We represent the region divided into thirds by intensity with Q1, Q2, and Q3, using 0.33 and 0.66 as thresholds. By observing the number of instances modeled, we observe that the spherical vectors representing each emotion concentrate on the unique styles frequently associated with each emotion. This pattern reflects how people tend to express certain emotion styles more regularly in reality. For instance, in the case of surprise, a higher pitch is more commonly expressed than a lower pitch.
We observe that the prosodic pattern based on emotional style and intensity reflects the characteristics of the VAD axis. Positive valence is associated with higher pitch and energy as the emotional intensity increases, whereas the duration tends to shorten. Similarly, positive arousal leads to increased patterns of pitch and energy change but typically decreases patterns of change in duration. By contrast, with positive dominance, the pitch and energy variation ranges become narrower, whereas the duration tends to be broader. For example, in the happy emotion, we can observe a shift in the range between Q1 and Q3 as dominance changes, leading to distinct variations in “I” and “V” styles. Similarly, the “VII” style in the sad emotion reflects negative arousal, with prosodic variation patterns changing with intensity. Compared to the “VIII” style, the difference in valence results in noticeable differences in the average prosodic patterns.
V-A2 Average Prosodic Variation Rates
To demonstrate the diverse modeling of emotions, we used a metric that quantifies the overall emotion averages for prosody, energy, and duration and the average prosodic variation rates for the corresponding emotion styles. This metric $p_{c}$ is calculated using the formula $p_{s}-p_{a}/p_{a}$, where $p_{s}$ represents the average prosodic feature value for a specific emotion style and $p_{a}$ denotes the overall average prosodic feature value for that emotion. We observed different combinations of prosodic variation rates $p_{c}$ of pitch, energy, and duration for each emotional style. For example, we can observe that the cell hue combinations in the average prosodic variation rates vary across different emotions, highlighting the diversity in emotional expression. In summary, the results are consistent with our assumptions about speech emotion analysis, underscoring the efficacy of our modeling in accounting for emotional style diversity.
V-B Model Performance
We conducted experiments including seen and unseen scenarios to evaluate EmoSphere++ and baseline models for non-parallel style transfer. We split our experiments into two categories: 1) seen non-parallel style transfer and 2) unseen non-parallel style transfer.
V-B1 Seen Non-Parallel Style Transfer
We first demonstrate the robustness of our proposed model in seen non-parallel style transfer, where a TTS system synthesizes a seen speaker. We select another fixed pair of reference signals from the test set of the same emotion and speaker. As shown in Table II, EmoSphere++ outperforms the previous methods in terms of both subjective and objective evaluations. In terms of naturalness and linguistic consistency, EmoSphere++ achieves the highest nMOS with a score of 3.92 and performed strongly in WER with a score of 15.52 compared to those of the baseline models. Regarding emotion and speaker style similarity, EmoSphere++ scores the highest overall eMOS of 3.86 and sMOS of 3.97. The objective results of the speaker metric SECS and emotion metrics ECA, SVAS, and EECS show that EmoSphere++ outperforms state-of-the-art models in transferring custom speech styles. The proposed method demonstrates superior performance in style transfer and quality compared to previous approaches.
V-B2 Unseen Non-Parallel Style Transfer
Subsequently, we explored the robustness of the proposed model in the case of unseen non-parallel style transfer. We set up zero-shot scenarios with unseen speakers for the ESD dataset and tested how the TTS model reproduces each speaker style when synthesizing different emotional phrases. Mellotron [44] and Mixedemotion [10] are not suitable for zero-shot scenarios because they rely on speaker lookup tables and are fine-tuned to a single speaker. As shown in Table III, the result indicates the drop in overall metrics, indicating that adapting to unseen speakers is more complex than adapting to unseen emotions. However, the most consistent performance in the various emotion zero-shot scenarios suggests that EmoSphere++ can adjust to a wide range of unseen emotions.
V-C Emotion Intensity Control
V-C1 Comparison With Control Methods
As a comparative study, we implemented two speaker and emotion attribute control methods (i.e., Matcha-TTS w/ Scaling Factor, Matcha-TTS w/ Relative Attributes), as described in Section IV-D. We evaluated the performance of these two methods in terms of performance and control. To demonstrate the ability of our model to control intensity, we synthesized speech with three different levels of emotion intensity (weak, medium, and strong). In the relative attribute model and EmoSphere++, we uniformly refer to values of 0.1 as weak, 0.5 as medium, and 0.9 as strong. The scaling factor cannot assign intensity values; therefore, we set the scalar factor to 1, 2, and 3 to represent weak, medium, and strong emotion intensities, as in the original setting.
Performance. We explored the expressiveness of our proposed model in non-parallel style transfer, where a TTS system synthesizes both the prosodic style of a reference signal and modeling emotion attribute. The results are compiled and presented in Table IV for easy comparison. Compared to transferring emotion from label-based emotion with relative attributes and reference-based emotion with a scaling factor, improved speech quality and expressiveness are achieved using EASV.
Control. We calculated the average pitch of the synthesized speech across all test sentences based on the intensity of each emotion. Fig. 5 shows the following: 1) The Matcha-TTS w/ Relative Attributes reflects the adjustment of emotional properties, but it often shows a limited adjustable range and tends to be reduced to a more uniform style. This outcome suggests subtle emotional nuances cannot be easily captured using only emotion labels. 2) The Matcha-TTS w/ Scaling Factor exhibits the most variation, but it often becomes unstable when adjusted on labels such as sad. Therefore, determining an appropriate scaling factor is difficult, and adjustments may lead to instability in audio quality. Conversely, the pitch tendency plot of EmoSphere++ reflects the intensity based on emotion, indicating that the proposed model synthesizes speech according to the given intensity scale.
V-C2 Intensity Control in the Zero-Shot Scenario
To demonstrate the ability to control emotional expression in the zero-shot scenario, we visualized the tendency of the pitch as shown in Fig. 6. We calculated the average pitch of the synthesized speech for the unseen speaker across all test sentences based on the intensity of each emotion. The pitch trend graph changes with intensity, reflecting the nature of the emotion. We observe a decrease in pitch for the sad emotion, whereas the pitch tends to increase as intensity rises for other emotions. This pattern indicates that the synthesized speech in EmoSphere++ can control the intensity of each emotion, even for the zero-shot scenario.
TABLE VI: Ablation study on VAD extractor for parallel emotion transfer on IEMOCAP datasets.
Method WERWhis ($\downarrow$) WERw2v ($\downarrow$) WERAVG ($\downarrow$) SECSR ($\uparrow$) SECSW ($\uparrow$) SECSAVG ($\uparrow$) $\textbf{RMSE}_{f0}$ ($\downarrow$) $\textbf{RMSE}_{period}$ ($\downarrow$) F1 V/UV ($\uparrow$) ECA ($\uparrow$) SVAS ($\uparrow$) EECS ($\uparrow$)
GT 13.70 28.24 20.97 0.7383 0.8506 0.7945 - - - 54.50 - 0.6444
BigVGAN 14.93 13.95 14.44 0.6449 0.8496 0.7473 3.02 0.3262 0.6070 53.51 0.9685 0.6338
w/ Real VAD Value 28.51 46.25 37.38 0.6153 0.8283 0.7218 20.17 0.4580 0.4971 40.34 0.8407 0.5213
w/ Predict VAD Value (Proposed) 27.07 46.53 36.80 0.6149 0.8336 0.7243 19.95 0.4587 0.4997 39.50 0.8505 0.5450
Figure 9: Pitch tendency track according to intensity for each emotion. Pitch values were calculated by averaging the synthesized speech for each intensity across all test sentences.
V-D Emotion Style Shift
V-D1 Visual Comparisons
We visualized the prosodic attributes of pitch related to the emotion intensity to gain an intuitive understanding of emotion style. To illustrate the variation patterns in emotion intensity with shifted emotion style, we visualized the pitch track changes for a sample, both from seen and unseen speakers. As analyzed in Section V-A, the prosodic pattern based on emotional intensity reflects the characteristics of the AVD axis. As shown in Fig. 7, we visualized the pitch contour of sad utterances with the most varied styles. For example, style vectors with positive V axes have higher average pitch values and reduced duration; positive A axes have a pitch that tends to increase the changing patterns. By contrast, duration decreases, and positive D axes have a narrow range in changing patterns with a broader duration. These results indicate that the proposed EASV is meaningfully characterized and enables emotion style and intensity controllable speech synthesis in both seen and unseen.
V-D2 Comparison of Emotional Consistency
Fig. 8 shows the resulting ECA for representative combinations of style shift. Across all emotions, we observe a similar emotional consistency when shifting to the representative style combinations as when maintaining the original style. These results indicate that the proposed model effectively adjusts emotion styles while maintaining emotional consistency across all style transformations. Therefore, as hypothesized in various psychological theories [27, 28], spherical vectors of emotion style can be characterized as derivatives of basic emotions.
V-E Ablation Study
V-E1 Impact of Disentangling Module
We conducted ablation studies to evaluate the impact of the disentangling module used in EmoSphere++ and present the findings in Table V. We used comparative subjective metrics (nMOS, sMOS, and sMOS) and objective metrics (SECS, ECA, and EECS) to assess the expressiveness and quality of the generated speech. The increase in quality and expressiveness demonstrates the importance of disentangling the module between emotion and speaker embeddings. We also determined that our proposed normalized orthogonality loss significantly increases the expressiveness of the unseen case, indicating that it has an important role in providing strong generalization performance for zero-shot scenarios.
V-E2 Effectiveness of Predicted versus Real VAD Values
To compare the impact of SER models that are trained with emotion dimension annotations, we conducted comparative experiments on the interactive emotional dyadic motion capture database (IEMOCAP) dataset [59], which includes emotional dimension annotations. As a training input for the emotion-adaptive coordinate transformation, we compared two features: real VAD values from human labeled annotations and predicted VAD values from the SER. Table VI shows that the objective metrics of prosodic expressiveness for both real and predicted VAD values are similar, suggesting that the predicted values perform comparably to manual annotations. Furthermore, we demonstrated the pitch tendency plot in Fig. 9, where the predicted VAD values reveal a consistent and precise pattern across all emotions. The results indicate that the VAD values predicted from the SER provide reliable emotion style and intensity modeling comparable to the real VAD values. These results indicate that SER-based VAD extraction offers a dependable alternative to manual annotations, maintaining similar performance while improving control over emotional nuances.
Figure 10: Evaluation process rates the discriminability of synthesized speech samples across different intensity levels, with $W$, $M$, and $S$ representing weak, medium, and strong intensities, respectively.
V-E3 Comparison of Coordinate Transformation
To compare effective modeling, we compared the mean-based method of the spherical emotion vector (SEV) [3] and EASV. For a fair comparison, we compared intensity accuracy using speech pairs of the same emotion and style from the same model, differing only in the coordinate transformation. We referred to values of W as weak, M as medium, and S as strong. As shown in Fig. 10, all types of emotion intensity pairs (W$<$M, M$<$S, and W$<$S) demonstrate consistently high average accuracy for individual emotions and overall. These results suggest that, while the mean-based approach of SEV captures the target emotion to a certain extent, it can be further improved by considering the distribution of other emotional categories, such as variance. Our proposed EASV method builds on the SEV, leading to a more comprehensive emotional representation. Furthermore, the distributions of the target and other emotional categories demonstrate the effectiveness of modeling style and intensity for each
VI Discussion
This study represents an initial attempt at modeling and synthesizing emotion styles and intensities for controllable emotional speech synthesis. Although we have demonstrated the effectiveness of our method, some related issues remain unresolved. We discuss these issues and aim to inspire future studies.
VI-A Limitations of Data Imbalance
As mentioned in Section V-A, the emotional styles modeled in each spherical vector exhibit unique characteristics that are confined to specific emotions. These results indicate that people tend to express particular emotional styles more frequently in reality; therefore, we focused on using only representative styles. This simplification may pose a limitation in capturing the full range of emotional nuances. However, researchers can address this issue by expanding the model to include more diverse datasets. Moreover, this serves as the initial attempt to model and synthesize emotion styles and intensities, demonstrating the potential of this approach.
VI-B Remaining Challenges of SER Model
As we summarized the SER models in Section III-A, we utilized the fine-tuned wav2vec 2.0 for an SER task. In the study [48], VAD predicted by the wav2vec 2.0-based SER model was significant. Hence, we utilized VAD pseudo-labels to avoid the inherent subjectivity and high costs associated with data collection. However, the approach relies on the performance of the trained SER model and similar biases and challenges as those encountered in SER [48, 68]. We expect the extended VAD predictor [69] will solve this problem. Although we only experimented with predicting VAD values here, we can extend this to combine real VAD values across different datasets.
VII Conclusion
In this paper, we presented EmoSphere++, an emotion-controllable zero-shot TTS model that can control emotional style and intensity to resemble natural human speech. We presented a comprehensive analysis and study that characterizes emotional states as derivatives of primary emotions using a dimensional model at the initial stage. We propose a novel EASV to model emotional style and intensity as derivatives of primary emotions. We also introduced a controllable and expressive speech synthesis framework that utilizes a multilevel style encoder with additional loss functions. The experimental results thoroughly analyze the components of our model and demonstrate its ability to effectively synthesize and control speech performance, even in zero-shot scenarios. We demonstrated that controlling the spherical vector along the VAD axis enables independent emotional style and intensity control. Through ablation studies, we also showed that the proposed EASV effectively synthesizes the complex nature of emotion. This article only focused on studying emotion-controllable TTS for a limited set of emotions. Our proposed spherical vector can enable complete emotion control in most existing emotional speech synthesis frameworks. Furthermore, we plan to expand our experiments to include emotional voice conversion in future studies.